A cylinder tank has a 20 foot diameter and is 15 feet tall. How amny cubic feet of water cant the tank hold
Found 2 solutions by mananth, rfer:
Answer by mananth(16949)

A cylinder tank has a 20 foot diameter and is 15 feet tall. How amny cubic feet of water cant the tank hold
volume of a cylinder is given by pi*r^2*h where r is radius and h = height
= 3.142 * 10^2 * 15
4713 cu. ft.

Answer by rfer(16322)

v=pir^2h
v=3.14*100*15
v=4710 cu ft